OpenStudy (anonymous):

Factor: x2 + 2x - 15 (x - 2)(x - 9) (x - 3)(x + 5) (x - 3)(x - 5) prime

OpenStudy (chriss):

the trick is to find 2 numbers with a product equal to the last term and a sum equal to the coefficient of the middle term

OpenStudy (hoblos):

x2 + 2x - 15 = x2 + 5x -3x -15 = x(x+5)-3(x+5) = (x+5)(x-3)